A tightly stretched string of length is fastened at both ends. The midpoint is taken to a height 

   transversely and the string is released from rest in this position. Find the displacement of any point 

  of the string at any subsequent time.


1
Expert's answer
2022-01-10T14:54:48-0500

When string is released, then it will vibrate. Let the equation which represents the displacement is given by, "y(x,t)= Asin(\omega t +kx+\phi)"

where "\omega" is angular velocity, k is wave number, "\phi" is initial phase.


According to the question, at end points x=0 and x=l, displacement is zero and maximum displacement is at mid point that is h.

So, "A=h"


Now, since string is held at end points so it represents half wavelength of the wave. Then "k=\frac{2\pi}{\lambda}, \lambda =l/2 \implies k=\frac{4\pi}{l}"

For phase, at x=0 and x=l, displacement is zero, so

"0= hsin(\phi)" and "0 = hsin(\frac{4π}{l}l+\phi)= hsin(4π+\phi)"

so, "\phi=0"


velocity:

"y'_t=\omega Acos(\omega t +kx)"

"y'_t(x,0)=\omega Acos( kx)=0\implies \omega =0"


then:

"y(x,t)= hsin(4\pi x/l)"